# On the distance between mean and geometric median in high dimensions

### v1

@@ 1 Introduction
The geometric median provides a robust alternative to the sample mean, mitigaging the impact of outliers. The improved robustness comes at the cost of a bias: in the infinite-data limit, the geometric median is generally distinct from the mean (i.e., expectation) of the data generating distribution, which in practice often is the learning target [^1] For special distributions like the Gaussian, geometric median and mean can coincide. . A method to balance robustness and bias is the median-of-means [Lugosi2019, Minsker2015] or “bucketing” approach [Guerraoui2024].
@@ 4 Discussion
The bias of the geometric median was recently highlighted in [Minsker2024GeometricMedian], where a constant upper bound on the Euclidean bias for a large class of distributions is derived. We complement this result by showing the Euclidean bias to be $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ for a smaller class.
From an estimation viewpoint, a small distance to the mean is desirable. From a robustness viewpoint it raises concern, as the mean is not robust against, say, contaminating a portion of the observations. A differentiation clarifies this: while $M$-dependence is compatible with certain cellwise contamination where indeed the geometric median is not always robust [RAYMAEKERS2024Cellwise], rowwise contaminated observations are typically not $M$-dependent and the geometric median is robust here [Minsker2015, schwank2025robustscorematching].
There is empirical evidence that the geometric median may similarly be close to the mean for sparsely interacting random fields [schwank2025robustscorematching].
Our results link to the classic paper by [Brown1983] showing that (sample) geometric median and mean are equivalent in terms of asymptotic variance for isotropic Gaussians as the ambient dimension approaches infinity.
Finally, we remark that the expansion in Theorem 2.6 always contains the term $\operatorname{E}\left[\,(X_{i}-\mu_{i})^{3}\,\right]$, i.e., the negative component skewness. This aligns with statistical intuition for the univariate median [Hippel2005].
@@ A.1 Proofs
When $p\geq M+2$, then $X_{1}$ and $X_{M+2}$ are independent, which ensures the distribution of $\mathbf{X}$ does not concentrate on a one-dimensional linear subspace. By [Milasevic1987], the geometric median is unique. ∎
@@
by $\ell_{1}$ versus $\ell_{2}$ norm equivalence. Consequently, by [Shorack2017, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
@@
The Hoeffding-type bound from [Janson2004PartlyDependentHoeffding, Thm. 2.1] builds on $\chi(\mathcal{I})$, the number of sets containing fully independent variables needed to cover $\mathcal{I}$. By $M$-dependence, we may cover $\{1,\ldots,p\}$ by $M+1$ sets of independent variables: consider the index sets $\mathcal{I}_{j}:=\{j+l(M+1):l=0,\ldots,\big{\lfloor}\frac{p-j}{M+1}\big{\rfloor}\}.$ for $j=1,\ldots,M+1$. Then $(A_{p,i})_{i\in\mathcal{I}_{j}}$ are independent for all fixed $j=1,\ldots,M+1$. Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [Janson2004PartlyDependentHoeffding]. Thus
@@
First, recall that for any nonnegative random variable $Y$ and $j>0$, we have $\operatorname{E}\left[\,Y^{j}\,\right]=\int_{0}^{\infty}jy^{j-1}\operatorname{P}\left(\,Y>y\,\right)\mathrm{d}y$ [DurrettProbabilityBook, Lemma 2.2.13].
@@
Since by Lemma A.7 we have $\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\to 1$ in probability as $p\to\infty$, also $(\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]})^{-1/2}\to 1$ by the continuos mapping theorem. To show that the random sequence’s expectations also converge, we appeal to a uniform integrability argument [DasGupta2008, Thm 6.2]. Indeed, by Theorem 2.3 it holds for example that
@@
In summary, $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$ when $p$ large enough. By [DasGupta2008, Thm 6.2], we conclude $III\to\operatorname{E}\left[\,N/2\,\right]$ as $p\to\infty$.
@@
If $q\geq 4$, the claim can be shown using Chebychev’s inequality. To allow $q>3$ in Theorem 2.6, we instead employ a truncation approach similar to [DurrettProbabilityBook, Sect. 2.2.3]. For $1\leq i\leq p$, define $A_{p,i}:=(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}$ and set $a_{p}:=\operatorname{E}\left[\,\sum_{i=1}^{p}A_{p,i}\,\right]$.

### v2

@@ 1 Introduction
The geometric median provides a robust alternative to the sample mean, mitigaging the impact of outliers. The improved robustness comes at the cost of a bias: in the infinite-data limit, the geometric median is generally distinct from the mean (i.e., expectation) of the data generating distribution, which in practice often is the learning target [^1] For special distributions like the Gaussian, geometric median and mean can coincide. . A method to balance robustness and bias is the median-of-means [8, 10] or “bucketing” approach [5].
@@ 4 Discussion
The bias of the geometric median was recently highlighted in [11], where a constant upper bound on the Euclidean bias for a large class of distributions is derived. We complement this result by showing the Euclidean bias to be $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ for a smaller class.
From an estimation viewpoint, a small distance to the mean is desirable. From a robustness viewpoint it raises concern, as the mean is not robust against, say, contaminating a portion of the observations. A differentiation clarifies this: while $M$-dependence is compatible with certain cellwise contamination where indeed the geometric median is not always robust [12], rowwise contaminated observations are typically not $M$-dependent and the geometric median is robust here [10, 13].
There is empirical evidence that the geometric median may similarly be close to the mean for sparsely interacting random fields [13].
Our results link to the classic paper by [1] showing that (sample) geometric median and mean are equivalent in terms of asymptotic variance for isotropic Gaussians as the ambient dimension approaches infinity.
Finally, we remark that the expansion in Theorem 2.6 always contains the term $\operatorname{E}\left[\,(X_{i}-\mu_{i})^{3}\,\right]$, i.e., the negative component skewness. This aligns with statistical intuition for the univariate median [6].
@@ A.1 Proofs
When $p\geq M+2$, then $X_{1}$ and $X_{M+2}$ are independent, which ensures the distribution of $\mathbf{X}$ does not concentrate on a one-dimensional linear subspace. By [9], the geometric median is unique. ∎
@@
by $\ell_{1}$ versus $\ell_{2}$ norm equivalence. Consequently, by [14, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
@@
The Hoeffding-type bound from [7, Thm. 2.1] builds on $\chi(\mathcal{I})$, the number of sets containing fully independent variables needed to cover $\mathcal{I}$. By $M$-dependence, we may cover $\{1,\ldots,p\}$ by $M+1$ sets of independent variables: consider the index sets $\mathcal{I}_{j}:=\{j+l(M+1):l=0,\ldots,\big\lfloor\frac{p-j}{M+1}\big\rfloor\}.$ for $j=1,\ldots,M+1$. Then $(A_{p,i})_{i\in\mathcal{I}_{j}}$ are independent for all fixed $j=1,\ldots,M+1$. Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [7]. Thus
@@
First, recall that for any nonnegative random variable $Y$ and $j>0$, we have $\operatorname{E}\left[\,Y^{j}\,\right]=\int_{0}^{\infty}jy^{j-1}\operatorname{P}\left(\,Y>y\,\right)\mathrm{d}y$ [4, Lemma 2.2.13].
@@
Since by Lemma A.7 we have $\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\to 1$ in probability as $p\to\infty$, also $(\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]})^{-1/2}\to 1$ by the continuos mapping theorem. To show that the random sequence’s expectations also converge, we appeal to a uniform integrability argument [3, Thm 6.2]. Indeed, by Theorem 2.3 it holds for example that
@@
In summary, $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$ when $p$ large enough. By [3, Thm 6.2], we conclude $III\to\operatorname{E}\left[\,N/2\,\right]$ as $p\to\infty$.
@@
If $q\geq 4$, the claim can be shown using Chebychev’s inequality. To allow $q>3$ in Theorem 2.6, we instead employ a truncation approach similar to [4, Sect. 2.2.3]. For $1\leq i\leq p$, define $A_{p,i}:=(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}$ and set $a_{p}:=\operatorname{E}\left[\,\sum_{i=1}^{p}A_{p,i}\,\right]$.